Electric potential field is given by V=xy²z².find electric firld at P(-1,2,1)

Grade:12

1 Answers

amansingh
6 years ago
Ex=−∂v/∂x=−∂(xy2z2)/∂x=−y2z2=−4 volt/m
Ey=−∂v/∂y=−∂(xy2z2)/∂y=−2xyz2=4 volt/m
Ez=−∂v/∂z=−∂(xy2z2)/∂z=−2xy2z=8 volt/m
Enet=Ex+Ey+Ez=−4+4+8=8 volt/m